Find the number of grams in 180 drams, where 1 grain is 1/60 of a dram. Need to find the formula.: Find the number of grams in 180 drams, where 1 grain is 1/60 of a dram. Need to find the formula.
Answer by checkley77(3624)

160(1/60)=160/60=2.667 or 2 2/3 grams.